OpenStudy (anonymous):

Triangle ABC has side lengths 7, 24, and 25. Do the side lengths form a Pythagorean triple? Explain.

OpenStudy (anonymous):

yes

OpenStudy (hoblos):

25^2 = 625 24^2+7^2 = 625 so 25^2 = 24^2 + 7^2 thus they form a Pythagorean triple

sam (.sam.):

A Pythagorean triple has three positive integers a, b, and c, such that a^2 + b^2 = c^2. In other words, a Pythagorean triple represents the lengths of the sides of a right triangle where all three sides have integer lengths.Such a triple is commonly written (a, b, c). Some well-known examples are (3, 4, 5) and (5, 12, 13). A primitive Pythagorean triple is one in which a, b and c are coprime (the greatest common divisor of a, b and c is 1).

OpenStudy (anonymous):

its a perfect square